# A Coalition Game for On-demand Multi-modal 3D Automated Delivery System

### v1

@@ 2.1.1 Machine Learning Methods
Reinforcement learning (RL) has found applications in fleet dispatching and management problems such as on-demand delivery and various variants of vehicle routing problems. For example, (Jahanshahi et al., 2022) proposed a meal delivery service as a Markov Decision Process (MDP) and employed Deep Q-Networks (DQN) to optimize courier assignment integrating unique approach of order rejection and repositioning of idle couriers. This work extensively studied the real-time optimization of meal delivery by making order acceptance and courier deployment decisions. Likewise, (Mehra et al., 2023) introduces DeliverAI, a multi-agent reinforcement learning system that allows food delivery networks to provide dynamic routing by providing a novel distributed path-sharing algorithm using Q-learning. However, no graph-based approach was considered, and not all customers will be met by saving on the cost of an order delay in the former. The training process becomes more complex in both cases as models require specialized tuning and can be time-consuming.
Despite their effort to model the complex dynamic of such a problem, PDPTW, as a vehicle routing problem, properly falls into the sequence decision-making optimization, such as encoder and decoder schemes of pointer networks, recurrent neural networks and transformers optimizing by an RL agent. These methods are still modelled as MDP, but they mostly use policy-based methods such as the REINFORCE algorithm or actor-critic architecture to evaluate an action using a combined value-based approach (Williams, 1992). The encoder will process the input data and map them into high-dimensional representation. Subsequently, the decoder outputs each action step by step by extracting information from the internal memory by the design and comparing it with a critical baseline. The first deep learning model for sequential decision-making solution of VRP is introduced by Nazari et al. (2018). Afterwards, this optimization category opened up many research directions to which many scholars have diversely contributed. For instance, Mao et al. (2023) suggested a transformer encoder and LSTM with an attention mechanism for the decoder to model pickup and delivery problems. They used an actor-critic architecture equipped with a Generalized Advantage Estimator that can balance the bias and variance of the policy gradient estimates to show significant improvements over the most competitive baseline on the real-world dataset. Furthermore, James et al. (2019) proposed a novel deep reinforcement learning-based neural pickup and delivery optimization strategy to develop vehicle routing with a pointer network. This work has considered network-based routing, which utilizes the entire graph, not just the customer node, for distance cost calculation. They also study dynamic scenarios where the arrival time of new requests is emulated with a Poisson process and available energy for recharging changes randomly. Similarly, this study proposed an electric vehicle energy consumption model that considers a variety of factors, such as vehicle power consumption model, road information, and adaptive routing in terms of velocity and node choice based on the time window and physical obstacle, respectively. Regarding the electric vehicle delivery, Liu et al. (2023) proposed a specific UAV delivery problem with recharging characterized by directional edges and stochastic edge costs subject to wind conditions. Fuertes et al. (2023), also utilized multiple UAVs to solve vehicle routing for every agent considering the clustering of the customers and a variant of the Orienteering Problem.
These works either focused on real-world constraints in applying sequential optimization or designed a new algorithm for an RL-enhanced encoder-decoder. Although each gives more importance to one of the urban vehicle delivery aspects, such as routing throughout the entire network, realistic assumption of battery consumption and electric vehicle specification, they ignored incorporating the urban physical structure and dynamic constraints like congestion and obstacle avoidance. As a result, in this research, the scope of the problem is to be solved by RL augmented with a transformer, and the problem framework is defined to encompass a network-based urban vehicle routing problem that considers the dynamic city delivery structure. Furthermore, in the link-level analysis of the network, the edge embedding has been done in Lei et al. (2022) and illustrated that the encoder eventually produces a better node and graph embedding in most instances; thus, the edge-insertion technique in the graph attention will be used as a hop neighbourhood based on spatial and temporal time windows to capture the coupling distribution of the location and demand arrival.
Numerous other works concentrate on the encoder and decoder design while adding complexities like time windows and capacity, for instance, taking into consideration the edge information in the graph structure and residual connections between layers, which is called graph attention networks (GAT) by Veličković et al. (2017). In addition, GAT has shown significant improvements due to its inherent architecture of sharing information through edges; thereby, the entire graph is not required. Besides, Gao et al. (2020) proposed a modified GAT-based network that learns to build better heuristics for VRP and VRP-with-time windows for up to 400 nodes. They utilized a decoder and trained the network with an actor-critic network as an incentive for lower travel costs. Lei et al. (2022) presented edge-encoded GAT with a skip connection (residual E-GAT) at the encoder and used a pointer-based decoder with a self-attention module. Their research shows that edge embedding-based message aggregation has superior graph topology representation power. Likewise, Drori et al. (2020) employed GAT to solve routing problems in linear time. The results showed that the framework can generalize from training on random graphs to running on random and real-world graphs of a different sort. Finally, Zhang et al. (2023b) provided multi-depot vehicle routing problems with soft time windows and encoder-decoder architecture, which employs a graph attention network to extract the complex spatial-temporal correlations within time windows. From these works, it can be concluded that using GAT will reduce the computational burden of the encoding process due to the sparse attention of edges rather than including every single edge in the network. In this regard, Fellek et al. (2023) proposed an attention-based DRL model that learns an expressive graph structure by fusing edge and node information. To that end, the message passing assists in aggregating node features based on data from customer node links. After processing the feature, the encoder produces improved node and graph embeddings. This work reported a close result to the former work of (Lei et al., 2022), which both embedded all the network in the graph. In conclusion, we aim to design a customized heterogeneous GAT as an attention mechanism for pickup and delivery problems, taking the time windows into account to improve the efficiency of the transformer.
2.1.2 Transformer-based Methods
A great deal of research has been done in favour of the transformer detail design to enhance the quality of the result in terms of time, generalization, and scalability. Namely, the work by Rabecq and Chevrier (2022) accounted for capacitated pickup and delivery with time windows problems (CPDPTW) paired with an insertion heuristic that reduces the complexity of the feasibility check from exponential to quadratic. They have used an integrated encoding for the problem features, capacity and time windows. Similarly, Soroka et al. (2023) created CPDPTW and multiple depots and applied a reinforcement learning algorithm based on the Joint Attention Model for Parallel Route-Construction (JAMPR). They take additional routes and vehicle encoders to enrich the context for solving the vehicle routing problem with time windows in the following original work of transformer-based using multi-head attention mechanism (Kool et al., 2018). To do such, JAMPR creates a hidden embedding for each constructed route by embedding nodes and vehicle characteristics of the corresponding route and combining them with the context by the complete representation of the state, allowing multiple routes to be constructed in parallel. Furthermore, Son et al. (2023) proposed a new transformer model, Equity-Transformer, for min-max pickup and delivery routing problems. Such a transformer design contains two key inductive biases: multi-agent positional encoding for order bias and context encoder for equity context. The former considers additional depot nodes mimicking the single agent starting building tours. Conversely, the latter considers crucial factors such as temporal tour length, the target tour length, and the desired number of cities to be visited, thereby enhancing the fairness of the generated tours. Eventually, although work contributed to the vehicle routing field through a novel design and implementation, they did not incorporate the real-world case study or electric vehicle utilization in the urban environment delivery context or cope with multi-agent RL that does not repeat the routing with dummy depot, instead take all agent features into account for matching.
Most existing RL-based techniques, however, can only tackle ordinary VRPs, but the cooperative VRP problem with structural dependency and vehicle cooperation necessitates unique modelling. Li et al. (2021), for example, suggested a heterogeneous attention-based network to tackle single-vehicle PDP; however, cooperative PDP still needs to be solved. In contrast, the suggested Multi-agent PDP could accurately predict pickup-delivery dependencies and vehicle collaborations. On the other hand, these studies by Santiyuda et al. (2024); Zhang et al. (2022); Zhang et al. (2023a) presented the CPDPTW problem by multi-agents which specialize in the assigning and routing of the vehicles simultaneously. Each has designed a unique transformer architecture to account for better input data handling and increased computational efficiency. For example, the encoding approach comprises a joint encoding scheme to encode the spatiotemporal information in the first two, and the latter uses a graph clustering method based on the hop-neighbourhood network. Nevertheless, the decoding step is distinguished by particular context information, making them promising regarding vehicle assignment and checking the compatibility of the vehicle query by each node. Although all of them are composed of paired learning following the (Li et al., 2021) for precedence constraint as well as capacity and time windows, none of them properly addressed more complex constraints and random events in the urban on-demand delivery such as uncertain traffic environments. Furthermore, in the link-level analysis of the network, the edge embedding has been done in Lei et al. (2022) and illustrated that the encoder eventually produces a better node and graph embedding in most instances; thus, the edge-insertion technique in the graph attention will be used as a hop neighbourhood based on spatial and temporal time windows to capture the coupling distribution of the location and demand arrival.
2.2 Vehicle Cooperative Game
In the context of vehicle routing games, core game theory has been applied to various collaborative transportation and delivery problems, particularly focusing on cost-sharing mechanisms. For instance, Osicka et al. (2020) investigated the vehicle routing and location-routing problems (LRP), where decisions regarding facility location and routing are optimized collectively rather than independently. The study demonstrates that while core allocations are not guaranteed in collaborative location-routing scenarios, they frequently exist in most cases. Cooperative strategies can be beneficial as long as the core is not empty. It also provides insights into how coalition formations impact the viability and stability of shared logistics operations. Note that the cooperative game can be computationally intractable if the number of agents increases as the sub-coalition grows $2^{N}$, where $N$ is the number of agents. Therefore, the cost-sharing allocation for each sub-group can be cumbersome. To this end, Alexander et al. (2014) developed a novel direct coalition induction algorithm (DCIA) method for constructing the characteristic function to ensure that cooperative arrangements first exist and are dynamically stable over time. Additionally, the iterative coalition induction algorithm (ICIA) is introduced to adjust routing plans dynamically, maintaining optimal cooperation throughout the different stages of the game with no coalition having an incentive to leave the grand coalition. Moreover, Zibaei et al. (2016) in the problem of multi-depot vehicle routing problem, incorporated several methods from cooperative game theory, including the Shapely value, the Core, and $\tau$ value to allocate the cost savings fairly among the depot owners using numerical examples, where the cost savings and synergy effects of various coalitions are calculated, showing that the cost can be fairly distributed using cooperative game theory techniques. For even more expensive due to the need to solve the NP-hard vehicle routing problem for every possible coalition in the traditional cooperative game theory, Mak et al. (2023) proposed a deep multi-agent reinforcement learning (MARL) model as a coalitional bargaining game. This allows agents to implicitly map coalition formation to collaboration, gain space, and collect fair outcomes without access to the full characteristic function. However, in our research, we aim to find an optimal coalition involving a limited number of agents. Each coalition’s characteristic function comes from a sub-optimal solution to the vehicle routing problem.
@@ 2.2 Vehicle Cooperative Game
The review summary is depicted in Table 1. According to the table, our research has brought several aspects that have yet to be addressed, such as considering the uncertain and stochastic nature of the delivery system and external factors impacting the level of service for a multi-modal fleet in urban real-world constraints. As a result, this study, to the best of the authors’ knowledge, is the first that considers the on-demand food delivery problem modelled by electric-capacitated pickup and delivery with time windows (E-CPDPTW) using dynamic transformer while considering uncertain changes in the urban environment as well as network-based route planning in the presence of wind and physical city structure.
@@
NP TRL EV CVRP PDPTW UI MA EE Stochasticity
(James et al., 2019) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Zhang et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$
(Zhang et al., 2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Elsayed and Mohamed, 2020) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Zhang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Lei et al., 2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Fuertes et al., 2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Santiyuda et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Our Study $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
NP: Network Planning, RL: Transformer Reinforcement Learning, DAM: Dynamic Attention Mechanism, GAT: Graph Attention Network, EV: Electric Vehicle, CVRP: Capacitated Vehicle Routing Problem, PDPTW: Pickup and Delivery with Time Windows, MM: Multi-modal, UI: Urban Infrastructure EE: Edge-embedding, MA: Multi-agent

### v2

@@ 2.1 CE-CPDPTW Approaches
Initially, RL has found applications in fleet dispatching and on-demand delivery and various variants of vehicle routing problems. For example, Jahanshahi et al. (2022) proposed a food delivery service as a Markov decision process (MDP) using deep Q-networks (DQN) to optimize courier assignment. Likewise, Mehra et al. (2023) introduced DeliverAI, a multi-agent reinforcement learning system that allows food delivery networks to provide dynamic routing by a novel distributed Q-learning path-sharing algorithm. However, the training process becomes more complex in both cases as models require specialized tuning and can be time-consuming due to limited buffer memory.
Deep reinforcement learning, in addition, provides powerful memory-based architectures to account for sequence-to-sequence model learning, such as node visit sequence. These models are based on an encoder-decoder policy network, and mostly use policy-based methods such as the REINFORCE algorithm or actor-critic architecture to evaluate an action using a combined value-based approach (Williams, 1992). The encoder processes the initial features and maps them to a high-dimensional representation. Subsequently, the decoder outputs each action step by step by extracting information from the internal memory by comparing it with a critical baseline. The first deep learning model for the sequential decision-making solution of VRP is introduced by Vinyals et al. (2015), a pointer network for the travelling salesperson problem (TSP). Following this work, Nazari et al. (2018) argued that the pointer network fails in static features decoding; instead, they utilize the recurrent neural network (RNN) decoder coupled with an attention mechanism. Nevertheless, a wide variety of research has used transformers for both encoder and decoder due to the attention architecture, which can be designed to learn the node relationships and enhance the solution quality in terms of time, generalization, and scalability following Kool et al. (2018). For instance, Li et al. (2021) suggested a heterogeneous attention-based network to learn the pickup and delivery relation using six types of heterogeneous attentions of different nodes. A more general notion of this notion can be found in Löwens et al. (2022), where heterogeneous attention is sparsified based on chains of precedence constraints. However, their encoder embedding included only node features. To this end, Liu et al. (2023) proposed a UAV-based pickup and delivery problem, where attention layers are characterized by dot-product stochastic edge features merging into the node embedding. They represented a real-world case study delivery subjected to wind, and a stochastic edge-enhanced technique accounts for non-Euclidean distance. Their model outperforms Li et al. (2021) and Kool et al. (2018), though their scenario involves one-to-one matching and without a capacity constraint.
In particular, taking the edge information into account in the graph structure and residual connections between encoder layers has been shown to be more effective in VRP problems by Lei et al. (2022) and Fellek et al. (2023). Their research shows that edge embedding-based message aggregation has superior graph topology representation power due to its inherent architecture of sharing information through edges. This mechanism can help refine edge embedding based on its connectivity in the graph. However, in these studies, the entire graph has been incorporated in the GAT by distance-based edge weight, and all customers ended up adding unnecessary embedding for faraway nodes. Specifically, this can be worse when a node contains spatial and temporal features. For example, if two nodes are in a close distance in the graph, but they have a large gap in their time window visit, distance-based message passing does not reflect their true relationship, as they can be prioritized in encoding over those nodes which have an actual closer time window. In this regard, Zhang et al. (2023b) has demonstrated that it is necessary to take advantage of graph neural networks to extract local spatial–temporal information to improve the solution. As a result, edge embeddings in the graph attention network will be used, and an adjacency-based masking for close spatial and temporal features is applied to capture the coupling distribution of the location and demand time window, especially when there are two modes with two different networks. Another aspect of the importance of edge-enhancement is to take the urban network into account, which is more similar to a road-style network, where not necessarily a straight line can be found to link the nodes. For example, there might be several intermediate nodes connecting a pickup to its delivery node, where they are mostly considered as a direct Euclidean distance in the literature. In fact, James et al. (2019) proposed pickup and delivery vehicle routing with a pointer network, which considers network-based routing utilizing the entire graph, not just the customer node, for the cost adjacency matrix. Likewise, our study incorporates road-based network routing, where a larger graph is used from which a candidate subset of nodes is selected, and their connection is computed by Dijkstra’s algorithm.
In addition, most of the above-mentioned studies were simulated for single-vehicle systems and did not consider time window constraints for the pickup and delivery problem. Addressing more complex constraints, Soroka et al. (2023) added more layers and created a multi-vehicle CPDPTW, and took an additional encoder for vehicles to enrich the initial feature embedding. Despite the novel design performing well on a large-scale network, their model objectives only minimize the routing cost. In contrast, Santiyuda et al. (2024), Zhang et al. (2022), and Zhang et al. (2023a) presented the CPDPTW problem by multi-agent routing, simultaneously minimizing travel time and the late time arrival violation. Each has designed a unique transformer architecture to better handle the node relational information. For example, the encoding approach comprises a joint encoding scheme to encode the spatiotemporal information in the first two, and the latter uses a graph clustering method based on the hop-neighbourhood network. Nevertheless, although each work gives more importance to one of the real-world constraints of vehicle delivery, electric vehicle specifications like battery consumption, and the urban-related issues, such as incorporating physical structure to avoid obstacles, are ignored. As a result, to address these challenges regarding CE-CPDPTW using the UAV-ADR system, the scope of our methodology is defined as follows.
This study integrates a novel customized graph attention network encoder and a transformer decoder architecture using an end-to-end approach to solve CE-CPDPTW and minimize the delivery travel time and time delay. To do so, RL agents build up a tour in a sequential step through node to vehicle assignment with the highest probability based on the reward feedback from the decoder. Moreover, to better distinguish node-edge relationships for either of the modes in initial graph encoding, a dual heterogeneous encoder is introduced to account for precedence constraints and, more importantly, customers’ spatial and temporal correlation, which focuses on coupling relation learning of location and time window. The flowchart of this methodology is shown in Figure 2.
Figure 2: The reinforcement learning framework. Agents take a sequence of actions based on the reward feedback from the environment. Besides, the graph embedding encoder extracts the spatiotemporal joint correlation from the initial graph network features.
2.2 Collaborative Approaches
The MARL collaborative framework for air mobility has been studied by Fernando et al. (2023) in a decentralized manner under partial observations coupled with a fleet re-balancing mask to maximize the delivery fulfillment. Although decentralized training leads to learning independent policies for agents, and does need a global communication layer, fleet coordination can be challenging for a very large-scale fleet. Additionally, due to the sparse nature of the reward in such problems, training a decentralized cooperative policy is difficult, as it does not perform competitively (Park et al., 2021). On the other hand, a variety of studies presented centralized cooperative MARL with a global coordination layer. For instance, Son et al. (2023) proposed a novel equity-transformer for min-max pickup and delivery routing problems, which contains two key inductive biases: multi-agent positional encoding for order bias and context encoder for equity context. The former considers additional depot nodes mimicking the single agent starting building tours. Conversely, the latter considers crucial factors such as temporal tour length, the target tour length, and the desired number of cities to be visited, thereby enhancing the fairness of the generated tours. Furthermore, Fuertes et al. (2023) studied a system for cooperative is split into two stages: initial planning and routing solving. The initial planning proposes a clustering strategy to assign initial regions to each UAV, and the routing solving considers the initial and shared regions assigned during the initial planning to maximize the team reward for the orienteering problem. Moreover, Zong et al. (2022) presented a cooperative MARL for minimizing the travel cost, and demonstrated a significant difference in performance for unbalanced node distribution in the case of halting time in the simultaneous decision-making to resolve a node conflict. Lastly, Zhang et al. (2022) considered multiple ways of making action decisions, such as assigning all delivery tasks to vehicles based on a preset fleet order or the least travelling time. However, none of the research above has simulated a non-homogeneous fleet; therefore, no impact of interactive gain and cost allocation for each vehicle. In other words, there were no case studies indicating that a heterogeneous fleet can be exploited to minimize the system-level and common objective, such as customer waiting time. In fact, they either performed toward agent-specific cost, like minimum travel time or did not evaluate an effective combination of the fleet and profit distribution given real-world scenarios.
@@ 2.2 Collaborative Approaches
Finally, a summary of the related studies is depicted in Table 1. According to the table, our research has brought several aspects that have yet to be addressed, such as considering the uncertain and stochastic nature of the delivery system and external factors impacting the level of service for a multi-modal fleet in urban real-world constraints. As a result, this study, to the best of the authors’ knowledge, is the first that considers the on-demand food delivery problem modelled by collaborative electric-capacitated pickup and delivery with time windows (CE-CPDPTW) in the presence of wind and physical city structure.
@@
H-GAT TRL EV CVRP PDPTW UI MA EE Stochasticity
James et al. (2019) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zong et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Lei et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fuertes et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Santiyuda et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
[…]
@@
H-GAT: Heterogeneous Graph Attention Network, TRL: Transformer Reinforcement Learning, EV: Electric Vehicle, CVRP: Collaborative Vehicle Routing Problem, PDPTW: Pickup and Delivery with Time Windows, UI: Urban Infrastructure EE: Edge-embedding, MA: Multi-agent
In the next part, fundamental principles of coalition game theory, as well as cost allocation mechanisms, and how they can be applied to our problem, are discussed.
2.2.1 Coalition Game
One of the popular ways to investigate such facets is the coalitional game with a group of cooperating players. Coalition game theory has been applied to various centralized collaborative transportation and delivery problems, particularly focusing on cost-sharing mechanisms. For instance, the vehicle routing and location-routing problems, (Osicka et al., 2020). The study demonstrates that while core allocations are not guaranteed in collaborative location-routing scenarios, they frequently exist in most cases. Cooperative strategies can be beneficial as long as the core is not empty. Moreover, the multi-depot vehicle routing problem (Zibaei et al., 2016) and shared charging station for electric vehicles (Wang et al., 2023b) have also incorporated cooperative game theory, including the Shapley value and core analysis, showing that the cost can be fairly distributed.
The coalition game theory examines how groups are formed so that no member is incentivized to leave and play individually. In such games, agents agree on a cooperative task, maximizing the common goal (Lui, 2010). A coalition game $(K,C(S))$ is defined by a set of players $K$, and $C$ is called the characteristic function, a real number that represents the cost resulting from a coalition $S\subset N$ in the game $(K,C(S))$, which is a group of cooperating players. If $S=K$, the $C(K)$ is the cost of forming the coalition of all users, known as the grand coalition, with the assumption of $C(\emptyset)=0$. In addition, there is a cost allocation vector called user cost, $\phi_{i}$, which is the portion of $C(S)$ received by a player $i$ in coalition $S$ for each agent in coalitions (Chalkiadakis et al., 2022).
The first goal is to identify which coalitions can be formed to yield less cost owing to cooperative behaviour. The second is to determine a fair cost allocation mechanism by which potential coalitions should divide the user payoff among their members. Coalition game has several subclasses, such as monotone game, super-additive game, and convex game. Among which, a super-additive game is always profitable for two groups of players to join forces (Chalkiadakis et al., 2022). A game is called super-additive if it is defined by Equation 1, in which agents can always work without interfering with one another in disjoint coalitions, and no loss is involved in merging coalitions.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C(S_{1}\cup S_{2})\leq C(S_{1})+C(S_{2})\quad\forall S_{1},S_{2}\subseteq K:S_{1}\cap S_{2}=\varnothing}$ (1)
Where $S_{1}$ and $S_{2}$ are sub-coalitions from the greatest coalition exist; all agents working the coalition, therefore, $K$ is the total number of agents. Specifically, the cooperative case of our problem is established such that coalitions can be formed in two separate groups of UAV and ADR and not jointly from both groups simultaneously. In other words, each distinct coalition can consist of only UAVs or ADRs in its group. It is assumed that either these vehicles are owned by different entities or a parent company is planning to rent a profitable combination of each fleet. Therefore, a super-additive game can find coalitions with no incentive to leave the grand coalition. For example, if the UAV and ADR fleet sets are shown by $S_{U}$ and $S_{A}$, respectively, the ground coalition is then $K=S_{A}\cup S_{U}$. Therefore, CE-CPDPTW can be formed as a super-additive game if Equation 1 holds for the given coalition sets and their cost. The analysis of such games can be twofold. First is stability, according to Roger et al. (1991), the core of a game is defined as all allocation costs, such that no coalition wants to deviate from the grand coalition. The core can be empty as well, indicating no preference over cooperation. For the non-empty core, there are two necessary conditions regarding cost allocations, efficiency and rationality, which are shown in Equations 2 and 3, respectively.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{j\in N}\phi_{j}=C(K)}$ (2)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{j\in S}\phi_{j}\leq C(S)\quad\forall S\subseteq K}$ (3)
Upon satisfying these conditions, the core property of the game shows stability; thus, it incentivizes the formation of a grand coalition. Moreover, Shapley value fairly distributes the cost $C(K)$ that could be obtained by the grand coalition (Roth, 1988), to the unique cost allocation vector as shown in Equation 4. In this equation, each agent receives a proportional amount to their average marginal contribution, averaging over all the different sequences according to which the grand coalition could be built up from the empty coalition.
$\phi_{i}(K,C)=\frac{1}{K!}\sum_{S\subseteq K\backslash\{i\}}|S|!(|K|-|S|-1)![C(S\cup\{i\})-C(S)]$ (4)
Therefore, for a given characteristic function and combination of disjoint groups, the Shapley value can be used to split revenue or savings, based on the cost function.